OpenStudy (anonymous):

find the derivative of f(x)=arcsin(4(lnx))

OpenStudy (anonymous):

\[f \left( x \right)=\sin^{-1} \left( 4\left( \ln x \right) \right),f'(x)=\frac{ 1 }{ \sqrt{1-16\left( \ln x \right)^2} }*\frac{ 4 }{x }\]

OpenStudy (anonymous):

Thank you !!! :)

OpenStudy (anonymous):

is it the correct answer???

OpenStudy (anonymous):

i don't know what the answer is, but calculation is correct.

OpenStudy (anonymous):

I think so too but some how it is showing me as it is wrong